OpenStudy (anonymous):

The equation of the circle whose center is (0, 0) and radius is length 9 is x² + y² = 3 x² + y² = 9 x² + y² = 81

OpenStudy (anonymous):

\[\huge x^2 + y^2 = 9^2 \rightarrow x^2 + y^2 = 81\] @gazed